OpenStudy (anonymous):

What is the plot and slope for y=-4x-3?

OpenStudy (anonymous):

What are you asking?

OpenStudy (anonymous):

what is the x and y axis of y=-4-3

OpenStudy (anonymous):

y = mx + b m= the slope y= -4x - 3 m= -4 (the slope)

OpenStudy (anonymous):

Find Y and X coordinates by plugging in 0 respectively. When x = 0, y = 3 (This is one point) When y = 0, x= (-3/4)